OpenStudy (anonymous):

9(-5/18)=? (Show Work)

OpenStudy (anonymous):

\[9=\frac{9}{1}\]\[\frac{9}{1}*\frac{-5}{18}=\frac{9*-5}{18}\] 18=9*2